20231212
Time
1600-???
Attendance
Onkar, Roy
Spaces Occupied
PHYS 401, 433c, Dome
Goals
- Practice taking calibration frames
- Practice setup/shutdown
- Practice blogging
- Collect frames for limiting mag study
- Calibration Frames
- Light Frames of star fields
Frame Table
Sources
Time Observed | Source Name | Filter | Int Time [s] | Mag (filter) | RA [h] | DEC [deg] | Source Type | Scope | Purpose | Notes |
---|---|---|---|---|---|---|---|---|---|---|
20231212T1180 | HD 198478 | V | 0.6 | 4.8(V) | 20.75 | 46 | Star | Main | Seeing | |
20231212T1822 | GD 391 | U | 25.7 | ~13(V) | 20h30m59.32s | 39d38'27.6 | Starfield | Main | Limiting Mag | Sb:0.83e/px, g150 |
| 20231212T1835 | GD 391 | B | 2.055 | ~13(V) | 20h30m59.32s | 39d38'27.6 | Starfield | Main | Limiting Mag | Sb:10.33e/px, g150 |
20231212T1805 | GD 391 | V | 1.715 | ~13(V) | 20h30m59.32s | 39d38'27.6 | Starfield|Main | Limiting Mag | Sb: 12.38e/px, g150 |
| 20231212T1840 | GD 391 | R | 0.895 | ~13(V) | 20h30m59.32s | 39d38'27.6 | Starfield | Main | Limiting Mag | Sb:23.73e/px, g150 |
20231212T1842 | GD 391 | I | 3.5 | ~13(V) | 20h30m59.32s | 39d38'27.6'' | Starfield | Main | Limiting Mag | Sb:6.06, g150 |
20231212T1858 | NGC 6946 | V | ~10(V) | 20.5h | 60d | GALAKSHE | Main | Limiting Mag |
Calibration Frames
Time Taken | Calib Type | Filter | Int Time [s] | Scope | Gain | Purpose | Notes |
---|---|---|---|---|---|---|---|
20231212T1522 | FLAT | U | 20x 0.08s | Main | 150 | Light Calib | NINA |
20231212T1522 | FLAT | B | 20x 0.02s | Main | 150 | Light Calib | NINA |
20231212T1522 | FLAT | V | 20x 0.02s | Main | 150 | Light Calib | NINA |
20231212T1522 | FLAT | R | 20x 0.02s | Main | 150 | Light Calib | NINA |
20231212T1522 | FLAT | I | 20x 0.14s | Main | 150 | Light Calib | NINA |
Log
- Normal startup procedure, mirror doors opened without issue
- Took sky flats at 1522 in all five filters. Saved in Z:ObservatoryData/NINA/2023-12-12
- Removed flats from 2023-12-11, since the IR lamp was on in the dome
- Pointed at Deneb and set object position approx 1630
- Calibrated phd2 on Deneb
- Went from Deneb to HD 198478 for seeing estimations
- SharpCap Measured 5.0 +- 0.1
- PHD2 Measured 4.5“. HFD 6.5
- Three frames collected and stored in Z:ObservatoryData/SharpCap/2023-12-12
ToDo
Notes
Weather
- 1730
- Dome: 10C
- Ambient: 9C
- Humidity 38%